# A Survey on Text-guided 3D Visual Grounding: Elements, Recent Advances, and Future Directions

### v1

@@ Abstract
Text-guided 3D visual grounding (T-3DVG), which aims to locate a specific object that semantically corresponds to a language query from a complicated 3D scene, has drawn increasing attention in the 3D research community over the past few years. Compared to 2D visual grounding, this task presents great potential and challenges due to its closer proximity to the real world and the complexity of data collection and 3D point cloud source processing. In this survey, we attempt to provide a comprehensive overview of the T-3DVG progress, including its fundamental elements, recent research advances, and future research directions. To the best of our knowledge, this is the first systematic survey on the T-3DVG task. Specifically, we first provide a general structure of the T-3DVG pipeline with detailed components in a tutorial style, presenting a complete background overview. Then, we summarize the existing T-3DVG approaches into different categories and analyze their strengths and weaknesses. We also present the benchmark datasets and evaluation metrics to assess their performances. Finally, we discuss the potential limitations of existing T-3DVG and share some insights on several promising research directions. The latest papers are continually collected at https://github.com/liudaizong/Awesome-3D-Visual-Grounding.
@@ 1 Introduction
Three-dimensional (3D) visual data has emerged as a pivotal component in daily life, owing to advancements in sensing technologies and computer vision [1, 2, 3]. Unlike traditional 2D images, 3D scenes are generally represented as more complicated unordered point clouds, capturing richer spatial and depth information and introducing unique challenges and complexities due to the increased dimensionality and the intricacies of geometric and semantic interpretations. Therefore, 3D scenes have natural advantages for multimedia intelligence exploration and research, drawing significant attention to related real-world applications such as robotic navigation [4], computer-aided room design [5], and human-computer interaction [6]. However, the raw 3D scenes are too redundant with multiple different types of objects, and of high information sparsity against the user-specific retrieval demands for finding one object. Furthermore, it is also challenging to maintain and manage these raw 3D scenes since they need to occupy a huge number of storage resources. Therefore, the ability to quickly retrieve a specific object from a 3D scene can allow users to locate highlighted objects of their interests conveniently and help information providers to optimize the storage fundamentally, thus being of great importance and interest in the research community.
1.1 Definition and History
@@
Given a point cloud based 3D scene, the text-guided 3D visual grounding (T-3DVG) task [7, 8] aims to locate a specific 3D object semantically according to a free-form language description. For example, as shown in Fig. 1, for the sentence description “This light brown couch is next to a tree. It is in front of a big window. It is next to a chair with a blue and light brown pattern”, T-3DVG is required to predict the precise bounding box (the green one in the figure) of the corresponding object “couch”.
Figure 1: An illustration of text-guided 3D visual grounding (T-3DVG).
@@
As a fundamental but challenging vision-language task, T-3DVG serves as an intermediate step for various downstream vision-language tasks, such as multi-modal 3D scene understanding [9, 10], language-guided robotic grasping [11], and 3D embodied interaction [12]. These tasks need to first localize relevant objects according to given queries (image or text), and then correlate the retrieved objects with the spatial-related backgrounds within the scene to reason and output suitable intermediate features, answers or actions. Therefore, T-3DVG serves a critical role and is important for providing robust reasoning ability for many existing 3D-related multi-modal tasks.
@@
T-3DVG also shares similarities with some classic multi-modal tasks. For instance, text-guided 2D video grounding [13, 14, 15] and image grounding [16, 17] aim to locate an activity-specific moment/object related to a sentence query from a video/image. These 2D tasks require matching the semantic features of video frames or images with the query to determine the moment boundary or object location, respectively. In contrast, T-3DVG solely has coordinate-wise points instead of prior knowledge of any instance to match with the sentence. Besides, T-3DVG is also more challenging as scene data contains both complex spatial relationships and complicated background contents compared to 2D images. Therefore, T-3DVG needs to comprehensively understand the spatial relations of different objects from the textual description and reason them in scene data. T-3DVG is also similar to image-guided 3D detection [18, 19, 20], which aims to locate a specific object in outdoor scenes by querying its 2D pictures. Since the same information on object appearance and shape is shared between images and point clouds, this task only needs a simple matching between cross-modal semantics. Different from it, T-3DVG is more challenging as its language guidance solely provides textual information but with complex spatial correlations, making it difficult to retrieve the objects in complicated scenes due to their huge modality gap.
@@
T-3DVG was proposed in year 2020 [7, 8]. The research works of text-guided 3D visual grounding published in the past few years can be roughly categorized into two types: two-stage and one-stage frameworks. Specifically, the two-stage framework [21, 22, 23] first defines multiple object proposals through a pre-trained 3D detector, then reasons the objects with the language text to determine the best one. Although this type of framework has achieved significant performance, it relies on the quality of the 3D detectors and requires time-consuming reasoning among the multiple proposals while the frameworks are not end-to-end. Without using object proposals, the one-stage framework [24, 25, 26] takes the whole 3D scene as the input and embeds point-level features in an end-to-end manner. After point-text cross-modal reasoning, it directly regresses the spatial bounding box of the text-guided object. However, this framework overlooks the rich information between the global points and fails to model the object-level correlations, thus their predicted object proposals obtained are relatively coarse. To address the above challenges, intensive works [27, 28] have been proposed to design better feature encoders and cross-modal interactors in the early stage. Later, additional contextual semantics (like 2D assistant data, multi-view data) [29, 30] are investigated to improve the robustness and generalization of the text-scene understanding. Different settings like fully- and weakly-supervised learning [7, 31] have also been explored by the researchers for implementing this task in real-world applications. Multiple types of architectures, such as CNN-based [7], MLP-based [8], and Transformer-based [27], are further introduced to tackle the task from different perspectives.
1.2 Motivation and Significance
@@
However, without detailed guidance, the availability of a large number of existing T-3DVG methods can easily confuse researchers or practitioners trying to select or design an algorithm suitable for the specific problem at hand. To this end, this paper presents a comprehensive taxonomy of the recent approaches and developments of the text-guided 3D visual grounding research. We collect papers from reputable conferences and journals in CV, NLP, MM, IR, and machine learning areas, e.g., CVPR, ECCV, ICCV, WACV, EMNLP, ACM MM, NeurIPS, AAAI, etc. The papers were mainly published from 2020 to 2024. Fig. 2 summarizes the statistics of the collected papers. By abstracting common generalities in all methods, we summarize the different types of 3D grounding methodologies and reveal a common pipeline of the T-3DVG model. We also give a detailed overview of the general elements in existing T-3DVG models and discuss their strengths and weaknesses. At last, we conclude more concrete and promising future research directions (such as using powerful large models, and designing more practical settings) in this survey paper.
@@
Our goal is to provide the above perspectives on the composition of T-3DVG methods and the recent state of T-3DVG research. With such perspectives, informed practitioners can confidently evaluate the trade-offs of various T-3DVG approaches and make informed decisions about using a suite of techniques to design a T-3DVG solution.
@@
The survey is organized as follows. In Section 2, we introduce the general pipeline of the T-3DVG method and illustrate the technical details in a tutorial manner. It provides the reader with background knowledge about the general appearance and functional components of T-3DVG models. Section 3 summarizes the basic benchmark datasets and evaluation metrics. Section 4 divides T-3DVG solutions into several categories, elaborates on the methods within each category, and discusses their advantages and disadvantages. Section 5 summarizes the current research progress through performance comparison. Section 6 discusses open issues and further research directions. Section 7 concludes the paper.
2 Background
@@
Currently, there is no universal framework or process for text-guided 3D visual grounding (T-3DVG), and various methods propose novel architectures and techniques to address this problem from different perspectives. However, we observe that they often share common modules, as shown in Fig. 3. Before introducing the existing methods of T-3DVG, this section provides the details of some general tools (i.e., feature extractor, feature encoder, multi-modal interaction module, grounding head) as preliminaries.
2.1 Basic Notations and Preliminary
@@
Before introducing each general module, we first define the basic notations involved in T-3DVG. T-3DVG typically takes inputs from two modalities. One is the scene point cloud $\mathcal{P}\in\mathbb{R}^{N\times(3+K)}$, where $N$ is the number of the points, and the feature dimension $3+K$ denotes three-dimensional coordinates and $K$-dimensional additional features (such as color, normals, height and multiview features). The other input is the descriptive text $\mathcal{T}=\{w_{n}\}_{n=1}^{L}$, where $w_{n}$ is the $n$-th word and $L$ is the number of the words. The task of T-3DVG aims to locate a specific object related to the semantic of $\mathcal{T}$ in the scene $\mathcal{P}$. The ground truth bounding box of the target object is denoted by $\mathcal{D}\in\mathbb{R}^{(6+18)}$, where the dimensions of $\mathcal{D}$ include 3D center point coordinates, 3D bounding box size, and an 18-dimensional one-hot label for the category (in some cases). Therefore, the goal of the T-3DVG can be formulated as: $\mathcal{F}_{T-3DVG}:(\mathcal{P},\mathcal{T})\xmapsto{}\mathcal{D}$.
2.2 3D Scene Feature Extractor
@@
Figure 3: A general pipeline for the T-3DVG task, consists of multi-modal feature extractors, multi-modal feature encoders, multi-modal interaction module, and the final grounding head. An additional object detector is further introduced for two-stage T-3DVG methods to pre-extract all possible object proposals in each scene.
@@
Scene preprocessor. Generally, the number of points contained within the 3D scene point cloud of the entire ScanNet dataset [32] can be extremely large and variable. Processing each point in the original point cloud is typically costly and not conducive to parallel computation. Specifically, each scene in the ScanNet dataset contains about 40 million points. Therefore, it is reasonable to downsample all of them to a certain number of points. The T-3DVG method typically begins by randomly sampling $N_{p}$ points from the scene point cloud $\mathcal{P}$ as input points $\mathcal{P^{\prime}}\in\mathbb{R}^{N_{p}\times(3+K)}$ ($N_{p}$ is commonly set to 40,000 in practice). Then, the point cloud undergoes traditional data augmentation strategies consisting of random flipping, random rotation, random translation, and random size scaling for robust learning.
Object-level feature extractor. After preprocessing the original scene data, the T-3DVG methods with a two-stage framework generally utilize a pre-trained 3D object detector or segmentor to extract candidate objects and encode them at the object level. These object-level features are utilized to interact with the sentence query for reasoning the best match object. Commonly used 3D object detectors include Votenet [33], GroupFree [34], PointGroup [35], etc.. Some other methods propose extracting instances from the point cloud using a segmentation model [36, 37, 38, 39, 40] to preserve more fine-grained features. Specifically, these two-stage T-3DVG methods typically detect all possible candidate objects in the point cloud and obtain their localization $\mathcal{D}\in\mathbb{R}^{M\times 6}$ and their features $\mathbf{C}\in\mathbb{R}^{M\times d}$, where $M$ is the number of candidate objects and $d$ is the dimension of the objects’ feature. The feature extraction process can be described as:
@@
The localization of each object $\mathcal{D}_{i}\in\mathbb{R}^{6}=(c_{x},c_{y},c_{z},r_{x},r_{y},r_{z})$ is represented by its 3D center coordinates $c$ and 3D box length $r$.
Point-level feature extractor. The T-3DVG method with a single-stage framework does not explicitly detect or segment objects. Instead, these methods directly utilize a point cloud encoder, usually employing PointNet++ [41] or 3D convolutions[42], to obtain point-level features, which are then used as extracted scene features. Compared to the object-level features, these point-level features are more fine-grained and can perceive more relevant background contexts. However, it also introduces much irrelevant background noise for ineffective reasoning. The point-level feature extraction process can be described as:
@@
2.3 Text Feature Extractor
Text preprocessor. The textual inputs $\mathcal{T}=\{w_{n}\}_{n=1}^{L_{w}}$ provided in the dataset consist of sentences of varying lengths. For preprocessing, each text query is typically segmented into tokens. Sentences longer than a certain threshold are truncated, and those shorter than the threshold are padded with zeros. Conventionally, the T-3DVG methods generally set this threshold to $L=126$. Subsequently, existing methods typically employ textual data augmentation by randomly replacing some tokens corresponding to object categories in sentences with the [UNK] token. Some methods employ additional data augmentation techniques, such as randomly masking out words [27] or combining multiple sentences within the same scene [43]. Some other methods [44] utilize large language models (LLM) to expand input text, thereby providing supplementary information, such as generating textual descriptions from different viewpoints compared to the original semantics.
@@ 2.3 Text Feature Extractor
2.4 Scene and Text Context Encoder
After separately extracting the rough features from the scene point cloud and text, many T-3DVG methods further design intra-modal encoders to capture the dependencies within each modality to obtain richer contextual information. Based on these more contextual features, they then map information from different modalities into the same space to facilitate subsequent interaction and matching. Despite using these contextual encoders, some early T-3DVG methods omit this encoding step and directly fuse the extracted rough features of the point cloud and text for subsequent processing.

### v2

@@ Abstract
Text-guided 3D visual grounding (T-3DVG), which aims to locate a specific object that semantically corresponds to a language query from a complicated 3D scene, has drawn increasing attention in the 3D research community over the past few years. Compared to 2D visual grounding, this task presents great potential and challenges due to its closer proximity to the real world, the complexity of data collection and 3D point cloud source processing. In this survey, we attempt to provide a comprehensive overview of the T-3DVG progress, including its fundamental elements, recent research advances, and future research directions. To the best of our knowledge, this is the first systematic survey on the T-3DVG task. Specifically, we first provide a general structure of the T-3DVG pipeline with detailed components in a tutorial style, presenting a complete background overview. Then, we summarize the existing T-3DVG approaches into different categories and analyze their strengths and weaknesses. We also present the benchmark datasets and evaluation metrics to assess their performances. Finally, we discuss the potential limitations of existing T-3DVG and share some insights on several promising research directions.
@@ I Introduction
Three-dimensional (3D) visual data has emerged as a pivotal component in daily life, thanks to advancements in sensing technologies and computer vision [1, 2, 3]. Unlike traditional 2D images, 3D scenes are generally represented as more complicated unordered point clouds, capturing richer spatial and depth information and introducing unique challenges and complexities due to the increased dimensionality and the intricacies of geometric and semantic interpretations. Therefore, 3D scenes have natural advantages for multimedia intelligence exploration and research, drawing significant attention to related real-world applications such as robotic navigation [4, 5], computer-aided room design [6, 7], and human-computer interaction [8, 9, 10]. However, raw 3D scenes are often too redundant with multiple different types of objects, and of high information sparsity against the user-specific retrieval demands for finding one object. Furthermore, it is also challenging to maintain and manage these raw 3D scenes since they need to occupy a huge number of storage resources. Therefore, the ability to quickly retrieve a specific object from a 3D scene can allow users to locate highlighted objects of their interests conveniently and help information providers to optimize the storage fundamentally, thus being of great importance and interest in the research community.
I-A Definition and History
@@
Given a point-cloud-based 3D scene, the text-guided 3D visual grounding (T-3DVG) task [11, 12] aims to locate a specific 3D object semantically according to a free-form language description. For example, as shown in Fig. 1, for the sentence description “This light brown couch is next to a tree. It is in front of a big window. It is next to a chair with a blue and light brown pattern”, T-3DVG is required to predict the precise bounding box (the green one in the figure) of the corresponding object “couch”.
Figure 1: An illustration of the text-guided 3D visual grounding (T-3DVG).
@@
As a fundamental yet challenging vision-language task, T-3DVG serves as an intermediate step for various downstream 3D multimedia tasks, such as multi-modal 3D scene understanding [13, 14], language-guided robotic grasping [15], and 3D embodied interaction [16]. These tasks need to first localize relevant objects according to given queries (image or text), and then correlate the retrieved objects with the spatial-related backgrounds of the scene to reason and output suitable intermediate features, answers, or actions. Therefore, T-3DVG plays a critical role in providing robust reasoning abilities for many existing 3D-related multi-modal tasks.
@@
T-3DVG also shares similarities with some classic multi-modal tasks. For instance, text-guided 2D video grounding [17, 18, 19] and image grounding [20, 21] aim to locate an activity-specific moment/object related to a sentence query from a video/image. These 2D tasks require matching the semantic features of video frames or images with the query to determine the moment boundary or object location, respectively. In contrast, T-3DVG solely has coordinate-wise points instead of prior knowledge of any instance to match the sentence. Besides, T-3DVG is also more challenging as scene data contains both complex spatial relationships and complicated background contents compared to 2D images. Therefore, T-3DVG needs to comprehensively understand the spatial relations of different objects from the textual description and reason them in scene data. T-3DVG is also similar to image-guided 3D detection [22, 23, 24, 25, 26], which aims to locate a specific object in outdoor scenes by querying its 2D pictures. Since the same information on object appearance and shape is shared between images and point clouds, this task only needs a simple matching between cross-modal semantics. In comparison, T-3DVG is more challenging as its language guidance solely provides textual information but with complex spatial correlations, making it difficult to retrieve objects in complicated scenes due to the huge modality gap between vision and language.
@@
T-3DVG was proposed in the year 2020 [11, 12]. The research works on this task published in the past few years can be roughly categorized into two types: two-stage and one-stage frameworks. Specifically, the two-stage framework [27, 28, 29] first defines multiple object proposals through a pre-trained 3D detector, then reasons the detected objects with the language text to determine the best one. Although this type of framework has achieved significant performance, it relies on the quality of the 3D detectors and requires time-consuming reasoning among the multiple proposals while the frameworks are not end-to-end trained. Without using object proposals, the one-stage framework [30, 31, 32] takes the whole 3D scene as the input and embeds point-level features in an end-to-end manner. After point-text cross-modal reasoning, it directly regresses the spatial bounding box of the text-guided object. However, this framework overlooks the rich information between the global points and sometimes fails to model the object-level correlations, thus their predicted object proposals are relatively coarse. To address the above challenges, intensive works [33, 34] have been proposed to design better feature encoders and cross-modal interactors in the early stage. Later, additional contextual semantics (like 2D assistant data, multi-view data) [35, 36] are investigated to improve the robustness and generalization of the text-scene understanding. Different settings like fully- and weakly-supervised learning [11, 37] have also been explored by researchers for implementing this task in real-world applications. Multiple types of architectures, such as CNN-based [11], MLP-based [12], and Transformer-based [33], are further introduced to tackle the task from different perspectives.
I-B Motivation and Significance
@@
However, without detailed guidance, the availability of a large number of existing T-3DVG methods can easily confuse researchers or practitioners trying to select or design an algorithm suitable for the specific problem at hand. To this end, this survey paper presents a comprehensive taxonomy of the recent approaches and developments of the text-guided 3D visual grounding research. We collect papers from reputable conferences and journals in Computer Vision, Natural Language Processing, Multimedia, and Machine Learning areas, e.g., CVPR, ECCV, ICCV, WACV, EMNLP, ACM MM, NeurIPS, AAAI, etc.. The papers were mainly published from 2020 to 2024. Fig. 2 summarizes the statistics of the collected papers. By abstracting common generalities in all methods, we summarize the different types of 3D grounding methodologies and reveal a common pipeline of the T-3DVG model. We also give a detailed overview of the general elements in existing T-3DVG models and discuss their strengths and weaknesses. At last, we conclude several concrete and promising future research directions (such as using powerful large models, and designing more practical settings) in this survey paper.
@@
Our goal is to provide the above perspectives on the composition of T-3DVG methods and the recent state of T-3DVG research. With such perspectives, researchers can not only confidently evaluate the trade-offs of various T-3DVG approaches in-depth and make informed decisions about using a suite of techniques to design a T-3DVG solution, but also appropriately implement suitable 3D grounding architecture in real-world applications.
@@
The survey is organized as follows. In Section II, we introduce the general pipeline of existing T-3DVG methods and illustrate the technical details in a tutorial manner. It provides the reader with background knowledge about the functional components of T-3DVG models. Section III divides T-3DVG solutions into several categories, elaborates on the methods within each category, and discusses their advantages and disadvantages. Section IV summarizes the basic benchmark datasets and evaluation metrics. Section V summarizes the current research progress through performance comparison. Section VI discusses open issues and further research directions. Section VII concludes the paper.
II Background
@@
Currently, there is no universal framework or process for text-guided 3D visual grounding (T-3DVG), and various methods propose novel architectures and techniques to address this problem from different perspectives. However, we observe that they often share common modules, as shown in Fig. 3. Before introducing the existing methods of T-3DVG, this section provides the general pipeline and details of some common tools (i.e., feature extractor, feature encoder, multi-modal interaction module, grounding head) as preliminaries.
II-A Basic Notations and Preliminary
@@
Before introducing each general module, we first define the basic notations involved in T-3DVG. T-3DVG typically takes inputs from two modalities. One is the scene point cloud $\mathcal{P}\in\mathbb{R}^{N\times(3+K)}$, where $N$ is the number of the points, and the feature dimension $3+K$ denotes three-dimensional coordinates and $K$-dimensional additional features (such as color, normals, height and multiview features). The other input is the descriptive text $\mathcal{T}=\{w_{n}\}_{n=1}^{L}$, where $w_{n}$ is the $n$-th word and $L$ is the number of the words. The task of T-3DVG aims to locate a specific object related to the semantic of $\mathcal{T}$ in the scene $\mathcal{P}$. The ground truth bounding box of the target object is denoted by $\mathcal{D}\in\mathbb{R}^{b}$, where the dimensions $b$ of $\mathcal{D}$ include a 6-dimensional 3D center point coordinates, 3D bounding box size, and an 18-dimensional one-hot label for the category (in some cases). Therefore, the goal of the T-3DVG can be formulated as: $\mathcal{F}_{T-3DVG}:(\mathcal{P},\mathcal{T})\xmapsto{}\mathcal{D}$.
II-B 3D Scene Feature Extractor
@@
Figure 3: A general pipeline for the T-3DVG task, consisting of multi-modal feature extractors, multi-modal feature encoders, multi-modal interaction module, and the final grounding head. An additional object detector is further introduced for two-stage T-3DVG methods to pre-extract all possible object proposals in each scene.
@@
Scene preprocessor. Generally, the number of points contained within the 3D scene point cloud (e.g., the commonly adopted ScanNet dataset [38]) can be extremely large and variable. Processing each point in the original point cloud is typically costly and not conducive to parallel computation. For example, each scene in the ScanNet dataset contains about 40 million points. Therefore, it is reasonable to downsample all of them to a certain number of points for efficient processing. Existing T-3DVG methods typically begin by randomly sampling $N_{p}$ points from the scene point cloud $\mathcal{P}$ as input points $\mathcal{P^{\prime}}\in\mathbb{R}^{N_{p}\times(3+K)}$ ($N_{p}$ is commonly set to 40,000 in practice). Then, the point cloud undergoes traditional data augmentation strategies consisting of random flipping, random rotation, random translation, and random size scaling for robust learning.
Object-level feature extractor. After preprocessing the original scene data, existing T-3DVG methods with a two-stage framework generally utilize a pre-trained 3D object detector or segmentor to extract candidate objects and encode them at the object level. These object-level features are utilized to interact with the sentence query for reasoning the best matched object. Commonly used 3D object detectors include Votenet [39], GroupFree [40], PointGroup [41], etc.. Some other methods propose extracting instances from the point cloud using a segmentation model [42, 43, 44, 45, 46] to obtain more fine-grained features. Specifically, these two-stage T-3DVG methods typically detect/segment all possible candidate objects in the point cloud and obtain their localization $\mathcal{D}\in\mathbb{R}^{M\times 6}$ and their features $\mathbf{C}_{i}\in\mathbb{R}^{M\times d}$, where $M$ is the number of candidate objects and $d$ is the dimension of the objects’ feature. The localization of each object $\mathcal{D}_{i}\in\mathbb{R}^{6}=(c_{x},c_{y},c_{z},r_{x},r_{y},r_{z})$ is represented by its 3D center coordinates $\mathbf{c}$ and 3D box length $\mathbf{r}$. Overall, the feature extraction process can be described as:
@@
Point-level feature extractor. Existing T-3DVG methods with a single-stage framework does not explicitly detect or segment objects. Instead, these methods directly utilize a global point cloud encoder, usually employing PointNet++ [47] or 3D convolutions [48], to obtain point-level features, which are then used as extracted scene features. Compared to the object-level features, these point-level features are more fine-grained and can perceive more relevant background contexts. However, it also introduces much irrelevant background noise for ineffective reasoning. The point-level feature extraction process can be described as:
@@
II-C Text Feature Extractor
Text preprocessor. The textual inputs $\mathcal{T}=\{w_{n}\}_{n=1}^{L_{w}}$ provided in the dataset consist of sentences of varying lengths. For preprocessing, each text query is typically segmented into tokens. Sentences longer than a certain threshold are truncated, and those shorter than the threshold are padded with zeros. Generally, existing T-3DVG methods set this threshold to the same number of $L$. Subsequently, existing methods typically employ textual data augmentation by randomly replacing some tokens corresponding to object categories in sentences with specific Unknown token [UNK]. Some methods employ additional data augmentation techniques, such as randomly masking out words [33] or combining multiple sentences within the same scene [49]. Some other methods [50] utilize large language models (LLM) to expand input text, thereby providing supplementary information, such as generating textual descriptions from different viewpoints compared to the original semantics.
@@ II-C Text Feature Extractor
II-D Scene and Text Context Encoder
After separately extracting rough features from the scene point cloud and text, many T-3DVG methods further design intra-modal encoders to capture the dependencies within each modality to obtain richer contextual information. Based on these more contextual features, in the latter process, they then map information from different modalities into the same space to facilitate subsequent interaction and matching. Despite using these contextual encoders, some early T-3DVG methods omit this encoding step and directly fuse the extracted rough features of the point cloud and text for subsequent processing.